# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 11.2.3 Working Mechanism of the Ethereum Virtual Machine (EVM)
The Ethereum Virtual Machine (EVM) is the decentralized computational environment that executes smart contracts on the Ethereum network[215]. Every node on the Ethereum blockchain runs its own copy of the EVM, ensuring that the execution of contracts is consistent and verifiable across all nodes.
@@ 11.3.1 Decentralized Finance (DeFi)
Smart contracts have revolutionized the financial industry through the creation of Decentralized Finance (DeFi) applications[216]. DeFi allows users to access traditional financial services like lending, borrowing, and trading without intermediaries like banks or brokers. All transactions are governed by smart contracts, ensuring transparency and security.
For example, platforms like Aave[217] and Compound[218] allow users to lend and borrow cryptocurrency using smart contracts. Users deposit funds into smart contract-based pools, and borrowers can take loans by locking collateral into the same system. Interest rates are determined algorithmically based on supply and demand, and all transactions are executed automatically by the contracts.
@@ 11.3.2 Decentralized Autonomous Organizations (DAOs)
Decentralized Autonomous Organizations (DAOs) are organizations that are governed by smart contracts rather than centralized entities[219]. In a DAO, decision-making is automated and decentralized. Members can vote on proposals, and the rules for decision-making are enforced by smart contracts.
For instance, a DAO might be used to manage a shared investment fund. Members contribute funds to the DAO, and smart contracts control how the funds are invested or distributed based on the outcome of votes. One well-known example of a DAO is the MakerDAO[220], which governs the DAI stablecoin.
@@ 11.3.3 Gaming and Non-Fungible Tokens (NFTs)
Smart contracts have also made a significant impact on the gaming industry through the creation of Non-Fungible Tokens (NFTs)[221]. NFTs represent unique digital assets, such as in-game items, artwork, or collectibles, that can be bought, sold, and traded on blockchain platforms.
Smart contracts power NFTs by managing their creation, transfer, and ownership. For example, in games like ”Axie Infinity,”[222] players can collect and battle unique creatures (represented as NFTs). Each creature is distinct, and its ownership is recorded on the blockchain, ensuring scarcity and authenticity.
The following is an example of a smart contract used to create NFTs using the ERC721 standard[223]:
@@ 12.1.1 The Intent and Design of The DAO
The DAO (Decentralized Autonomous Organization) was one of the most ambitious projects launched on the Ethereum platform in 2016[224]. Its original intent was to create a decentralized investment fund that would operate entirely through smart contracts, without any intermediaries. The vision was to allow investors to pool their funds into The DAO, and vote on which projects should receive funding. This organization was revolutionary because it relied entirely on code to manage the fund’s governance, removing the need for a centralized authority[225].
The design of The DAO was simple in concept but complex in implementation. Participants could buy DAO tokens, which represented voting power in the organization[226]. Every decision was to be made through token-holder voting, and smart contracts would automatically execute the decisions based on the outcome of these votes. For example, if a project proposal was approved by a majority vote, funds would automatically be sent to the project team, all without human intervention.
One of the key advantages of The DAO was transparency and trustlessness[227] — since all of its operations were governed by code on the Ethereum blockchain, anyone could inspect the smart contract to verify how funds would be managed. In theory, this was a major step forward in decentralized governance, offering the possibility of creating a new way to fund and govern projects democratically.
@@ 12.1.2 Vulnerability Exploitation and Hacker Attack
Despite its ambitious design, The DAO’s code had a critical vulnerability. A flaw in the smart contract logic allowed an attacker to exploit a reentrancy bug[228], which enabled them to repeatedly withdraw funds from The DAO without properly updating the balance. This exploit worked by making recursive calls to the smart contract, draining funds each time before the contract could update its records to reflect the withdrawal.
The hacker was able to siphon off approximately 3.6 million Ether, worth about \$60 million at the time[229]. Here is a simplified explanation of the vulnerability:
@@ 12.1.3 Hard Fork: Ethereum and Ethereum Classic
In the aftermath of the hack, the Ethereum community faced a difficult decision. To address the situation and return the stolen funds, Ethereum developers proposed a hard fork[224]. A hard fork is a radical change to the blockchain protocol that results in the creation of two separate chains. In this case, the hard fork would effectively reverse the transactions made by the hacker, restoring the stolen Ether to its original owners.
@@
Ethereum 2.0, also known as Eth2 or ”The Merge,” is a major upgrade aimed at improving the scalability, security, and sustainability of the Ethereum network[230]. The current Ethereum blockchain, based on Proof of Work (PoW), has faced limitations such as high energy consumption, slow transaction speeds, and network congestion. Ethereum 2.0 seeks to address these issues through several key improvements:
@@ 12.2.2 Transition from PoW to PoS
The most significant change in Ethereum 2.0 is the transition from Proof of Work (PoW) to Proof of Stake (PoS) as the consensus mechanism[230]. In PoW, miners use computational power to solve complex mathematical puzzles and validate transactions. However, this approach is highly energy-intensive, leading to criticism about its environmental impact.
Proof of Stake (PoS) offers an alternative that requires far less energy. In PoS, instead of miners, there are validators who are chosen to create new blocks and validate transactions based on the amount of cryptocurrency they have staked as collateral[231]. The more tokens a validator stakes, the higher the chance they will be selected to validate the next block. This mechanism reduces the need for computational power and lowers energy consumption.
@@ 12.2.3 Sharding and Scalability
A key innovation of Ethereum 2.0 is the introduction of sharding, a technique designed to improve scalability[232]. Sharding divides the blockchain into smaller pieces, called shards, which can process transactions and smart contracts in parallel. This allows the network to handle a significantly larger volume of transactions, improving overall throughput.
@@ 12.3 Scalability Challenges of Ethereum
Ethereum, as one of the most popular blockchain platforms, faces significant scalability challenges[233]. As the number of users and decentralized applications (dApps) increases, the network’s limitations become more evident. These challenges can lead to high gas fees, slow transaction times, and network congestion. In this section, we will explore the key scalability issues Ethereum faces and examine some of the proposed solutions to improve its performance.
@@ 12.3.1 High Gas Fees
Gas in Ethereum measures the computational effort required to execute a transaction or smart contract[234]. Every operation in the Ethereum Virtual Machine (EVM), from basic arithmetic to storing data on the blockchain, has a specific gas cost. The total gas cost for a transaction is determined by summing the costs of all the operations required to process it.
When a user initiates a transaction, they specify the amount of gas they are willing to pay for its execution[235]. Miners prioritize transactions that offer higher gas fees since it is more profitable for them. This market-driven mechanism ensures that network resources are allocated efficiently, with users paying more during periods of high demand to have their transactions processed faster.
@@
  • Network Congestion: During times of high activity, such as popular dApp launches or NFT drops, the Ethereum network can become congested. With more users competing to have their transactions processed, gas prices rise as people are willing to pay more to get their transactions included in a block[236].
  • High Demand for dApps: Decentralized applications (dApps) like decentralized finance (DeFi) platforms and non-fungible token (NFT) marketplaces place significant demand on the network. Popular platforms like Uniswap[237] or OpenSea[238] can cause a spike in gas fees, as they require multiple complex smart contract interactions that consume a lot of gas.
@@
  • Batching Transactions: Batching allows users to group multiple transactions into one, saving on the total gas costs[239]. This is especially useful for frequent users or dApps that perform many small transactions.
  • EIP-1559: One of the most significant proposals to address high gas fees is Ethereum Improvement Proposal (EIP) 1559[240]. This proposal introduces a base fee that is burned, reducing the total supply of Ether over time, and a tip that incentivizes miners to include transactions. EIP-1559 also makes gas fees more predictable, reducing the chances of extreme price spikes during network congestion.
The ongoing development of Ethereum 2.0, which will introduce a proof-of-stake (PoS) consensus mechanism and shard chains, is expected to further alleviate gas fee issues by increasing the network’s capacity and throughput[241].
@@ 12.3.2 Exploration of Layer 2 Solutions
Layer 2 solutions are scaling methods that operate on top of the Ethereum main chain[242]. They help offload transaction processing from the main chain, reducing congestion and lowering gas fees. In this section, we will explore several Layer 2 technologies[242], including Rollups, state channels, and sidechains, which are designed to improve Ethereum’s scalability.
@@
Rollups are one of the most promising Layer 2 solutions for Ethereum scalability[243]. They allow transactions to be processed off-chain and then submit only compressed data or proofs to the Ethereum main chain. This significantly reduces the amount of data stored on-chain and increases transaction throughput.
@@
  • ZK-Rollups (Zero-Knowledge Rollups): ZK-Rollups use cryptographic proofs (called zero-knowledge proofs) to validate transactions[245]. These proofs ensure that the off-chain transactions are valid without the need to trust any third party. ZK-Rollups are generally faster and more efficient than Optimistic Rollups but are more complex to implement.
Rollups have gained significant traction in the Ethereum ecosystem due to their ability to reduce gas fees and increase transaction throughput while maintaining security. For example, Arbitrum[246] and Optimism[247] are leading platforms that use Optimistic Rollups, while zkSync[248] and StarkWare[249] leverage ZK-Rollups.
@@
State channels allow participants to conduct multiple off-chain transactions, with only the final state being committed to the blockchain[250]. This method is highly efficient for repeated interactions, such as payments or gaming applications, where the parties involved can settle their accounts off-chain and only submit the final balance to Ethereum.
@@
Plasma is a framework that allows for scalable and secure off-chain transactions[251]. Plasma chains are separate blockchains that run alongside the Ethereum main chain. They handle large volumes of transactions off-chain and interact with the main chain for security and settlement. Plasma chains can process thousands of transactions per second while ensuring that users can always retrieve their funds from the main chain if something goes wrong with the Plasma chain.
Sidechains[252], like Plasma, operate independently of the Ethereum main chain but still interact with it for finality and security. Sidechains have their own consensus mechanisms and can offer faster transaction times and lower fees. A popular example of a sidechain is Polygon (formerly Matic), which has become one of the leading Layer 2 solutions by providing fast and low-cost transactions.
@@ 12.3.3 Network Load from Decentralized Applications (dApps)
dApps can suffer from the same scalability issues as the Ethereum network itself, including high gas fees and slow transaction times. Popular dApps, particularly those in the DeFi space, often experience congestion during peak periods, making it difficult for users to interact with them affordably. For instance, during the DeFi summer of 2020, gas fees skyrocketed as users rushed to participate in yield farming[253], making smaller transactions prohibitively expensive.
@@
High-frequency activities, such as algorithmic trading on decentralized exchanges (DEXs) or frequent NFT minting, can place additional strain on the network[254]. These types of activities require a high number of transactions in a short period, leading to network congestion and increased gas prices.
@@
To overcome scalability challenges, many dApps are beginning to integrate Layer 2 solutions such as Rollups, state channels, and Plasma. For example, Uniswap[237], a leading decentralized exchange, has integrated with Optimistic Rollups to offer faster and cheaper trades. Similarly, gaming dApps are using state channels to allow users to interact with the platform off-chain, reducing the cost of frequent in-game transactions.
@@ 13.1.1 Definition of Proof of Stake
Proof of Stake (PoS) is a consensus mechanism that is utilized in blockchain networks to determine how the next block in a blockchain is chosen and validated[231]. Unlike Proof of Work (PoW), where participants (miners) solve complex cryptographic puzzles to create new blocks, PoS selects validators based on the amount of cryptocurrency they have ”staked” or locked up in the network[255].
In PoS, the idea is that those who have a larger stake in the network have more to lose if the system is compromised[255]. Hence, validators are selected to propose and validate new blocks in a way proportional to the amount of cryptocurrency they hold and are willing to stake. Validators are chosen through a pseudo-random selection process, with factors such as the size of the stake and the validator’s age (the time they have been staking) playing a role.
This method eliminates the need for intensive computational work, which is one of the main criticisms of PoW[256]. Instead, PoS is considered energy-efficient, as it removes the requirement for high energy consumption seen in mining operations.

### v3

@@ 11.2.3 Working Mechanism of the Ethereum Virtual Machine (EVM)
The Ethereum Virtual Machine (EVM) is the decentralized computational environment that executes smart contracts on the Ethereum network[ethereumEthereumVirtual]. Every node on the Ethereum blockchain runs its own copy of the EVM, ensuring that the execution of contracts is consistent and verifiable across all nodes.
@@ 11.3.1 Decentralized Finance (DeFi)
Smart contracts have revolutionized the financial industry through the creation of Decentralized Finance (DeFi) applications[Jensen2021AnIT]. DeFi allows users to access traditional financial services like lending, borrowing, and trading without intermediaries like banks or brokers. All transactions are governed by smart contracts, ensuring transparency and security.
For example, platforms like Aave[aaveAave] and Compound[compoundv2CompoundUnderstanding] allow users to lend and borrow cryptocurrency using smart contracts. Users deposit funds into smart contract-based pools, and borrowers can take loans by locking collateral into the same system. Interest rates are determined algorithmically based on supply and demand, and all transactions are executed automatically by the contracts.
@@ 11.3.2 Decentralized Autonomous Organizations (DAOs)
Decentralized Autonomous Organizations (DAOs) are organizations that are governed by smart contracts rather than centralized entities[Ding2023ASO]. In a DAO, decision-making is automated and decentralized. Members can vote on proposals, and the rules for decision-making are enforced by smart contracts.
For instance, a DAO might be used to manage a shared investment fund. Members contribute funds to the DAO, and smart contracts control how the funds are invested or distributed based on the outcome of votes. One well-known example of a DAO is the MakerDAO[makerdaoMakerDAOUnbiased], which governs the DAI stablecoin.
@@ 11.3.3 Gaming and Non-Fungible Tokens (NFTs)
Smart contracts have also made a significant impact on the gaming industry through the creation of Non-Fungible Tokens (NFTs)[Arora2021SmartCA]. NFTs represent unique digital assets, such as in-game items, artwork, or collectibles, that can be bought, sold, and traded on blockchain platforms.
Smart contracts power NFTs by managing their creation, transfer, and ownership. For example, in games like ”Axie Infinity,”[axieinfinityAxieInfinity] players can collect and battle unique creatures (represented as NFTs). Each creature is distinct, and its ownership is recorded on the blockchain, ensuring scarcity and authenticity.
The following is an example of a smart contract used to create NFTs using the ERC721 standard[ethereumERC721NonFungible]:
@@ 12.1.1 The Intent and Design of The DAO
The DAO (Decentralized Autonomous Organization) was one of the most ambitious projects launched on the Ethereum platform in 2016[Mehar2017UnderstandingAR]. Its original intent was to create a decentralized investment fund that would operate entirely through smart contracts, without any intermediaries. The vision was to allow investors to pool their funds into The DAO, and vote on which projects should receive funding. This organization was revolutionary because it relied entirely on code to manage the fund’s governance, removing the need for a centralized authority[DuPont2017ExperimentsIA].
The design of The DAO was simple in concept but complex in implementation. Participants could buy DAO tokens, which represented voting power in the organization[PeaCalvin2024ConcentrationOP]. Every decision was to be made through token-holder voting, and smart contracts would automatically execute the decisions based on the outcome of these votes. For example, if a project proposal was approved by a majority vote, funds would automatically be sent to the project team, all without human intervention.
One of the key advantages of The DAO was transparency and trustlessness[Kondova2019GovernanceOD] — since all of its operations were governed by code on the Ethereum blockchain, anyone could inspect the smart contract to verify how funds would be managed. In theory, this was a major step forward in decentralized governance, offering the possibility of creating a new way to fund and govern projects democratically.
@@ 12.1.2 Vulnerability Exploitation and Hacker Attack
Despite its ambitious design, The DAO’s code had a critical vulnerability. A flaw in the smart contract logic allowed an attacker to exploit a reentrancy bug[Liu2018ReGuardFR], which enabled them to repeatedly withdraw funds from The DAO without properly updating the balance. This exploit worked by making recursive calls to the smart contract, draining funds each time before the contract could update its records to reflect the withdrawal.
The hacker was able to siphon off approximately 3.6 million Ether, worth about \$60 million at the time[Dingman2019DefectsAV]. Here is a simplified explanation of the vulnerability:
@@ 12.1.3 Hard Fork: Ethereum and Ethereum Classic
In the aftermath of the hack, the Ethereum community faced a difficult decision. To address the situation and return the stolen funds, Ethereum developers proposed a hard fork[Mehar2017UnderstandingAR]. A hard fork is a radical change to the blockchain protocol that results in the creation of two separate chains. In this case, the hard fork would effectively reverse the transactions made by the hacker, restoring the stolen Ether to its original owners.
@@
Ethereum 2.0, also known as Eth2 or ”The Merge,” is a major upgrade aimed at improving the scalability, security, and sustainability of the Ethereum network[Asif2023ShapingTF]. The current Ethereum blockchain, based on Proof of Work (PoW), has faced limitations such as high energy consumption, slow transaction speeds, and network congestion. Ethereum 2.0 seeks to address these issues through several key improvements:
@@ 12.2.2 Transition from PoW to PoS
The most significant change in Ethereum 2.0 is the transition from Proof of Work (PoW) to Proof of Stake (PoS) as the consensus mechanism[Asif2023ShapingTF]. In PoW, miners use computational power to solve complex mathematical puzzles and validate transactions. However, this approach is highly energy-intensive, leading to criticism about its environmental impact.
Proof of Stake (PoS) offers an alternative that requires far less energy. In PoS, instead of miners, there are validators who are chosen to create new blocks and validate transactions based on the amount of cryptocurrency they have staked as collateral[Wendl2022TheEI]. The more tokens a validator stakes, the higher the chance they will be selected to validate the next block. This mechanism reduces the need for computational power and lowers energy consumption.
@@ 12.2.3 Sharding and Scalability
A key innovation of Ethereum 2.0 is the introduction of sharding, a technique designed to improve scalability[CortesGoicoechea2020ResourceAO]. Sharding divides the blockchain into smaller pieces, called shards, which can process transactions and smart contracts in parallel. This allows the network to handle a significantly larger volume of transactions, improving overall throughput.
@@ 12.3 Scalability Challenges of Ethereum
Ethereum, as one of the most popular blockchain platforms, faces significant scalability challenges[Bez2019TheSC]. As the number of users and decentralized applications (dApps) increases, the network’s limitations become more evident. These challenges can lead to high gas fees, slow transaction times, and network congestion. In this section, we will explore the key scalability issues Ethereum faces and examine some of the proposed solutions to improve its performance.
@@ 12.3.1 High Gas Fees
Gas in Ethereum measures the computational effort required to execute a transaction or smart contract[Khan2021GasCA]. Every operation in the Ethereum Virtual Machine (EVM), from basic arithmetic to storing data on the blockchain, has a specific gas cost. The total gas cost for a transaction is determined by summing the costs of all the operations required to process it.
When a user initiates a transaction, they specify the amount of gas they are willing to pay for its execution[Fajge2021WaitOR]. Miners prioritize transactions that offer higher gas fees since it is more profitable for them. This market-driven mechanism ensures that network resources are allocated efficiently, with users paying more during periods of high demand to have their transactions processed faster.
@@
  • Network Congestion: During times of high activity, such as popular dApp launches or NFT drops, the Ethereum network can become congested. With more users competing to have their transactions processed, gas prices rise as people are willing to pay more to get their transactions included in a block[Ko2022AnAO].
  • High Demand for dApps: Decentralized applications (dApps) like decentralized finance (DeFi) platforms and non-fungible token (NFT) marketplaces place significant demand on the network. Popular platforms like Uniswap[uniswapUniswapInterface] or OpenSea[opensea] can cause a spike in gas fees, as they require multiple complex smart contract interactions that consume a lot of gas.
@@
  • Batching Transactions: Batching allows users to group multiple transactions into one, saving on the total gas costs[Wang2021iBatchSE]. This is especially useful for frequent users or dApps that perform many small transactions.
  • EIP-1559: One of the most significant proposals to address high gas fees is Ethereum Improvement Proposal (EIP) 1559[ethereumEIP1559Market]. This proposal introduces a base fee that is burned, reducing the total supply of Ether over time, and a tip that incentivizes miners to include transactions. EIP-1559 also makes gas fees more predictable, reducing the chances of extreme price spikes during network congestion.
The ongoing development of Ethereum 2.0, which will introduce a proof-of-stake (PoS) consensus mechanism and shard chains, is expected to further alleviate gas fee issues by increasing the network’s capacity and throughput[Chen2023ComparisonOP].
@@ 12.3.2 Exploration of Layer 2 Solutions
Layer 2 solutions are scaling methods that operate on top of the Ethereum main chain[Neiheiser2023PracticalLO]. They help offload transaction processing from the main chain, reducing congestion and lowering gas fees. In this section, we will explore several Layer 2 technologies[Neiheiser2023PracticalLO], including Rollups, state channels, and sidechains, which are designed to improve Ethereum’s scalability.
@@
Rollups are one of the most promising Layer 2 solutions for Ethereum scalability[Thibault2022BlockchainSU]. They allow transactions to be processed off-chain and then submit only compressed data or proofs to the Ethereum main chain. This significantly reduces the amount of data stored on-chain and increases transaction throughput.
@@
  • ZK-Rollups (Zero-Knowledge Rollups): ZK-Rollups use cryptographic proofs (called zero-knowledge proofs) to validate transactions[Fernando2023PosterWA]. These proofs ensure that the off-chain transactions are valid without the need to trust any third party. ZK-Rollups are generally faster and more efficient than Optimistic Rollups but are more complex to implement.
Rollups have gained significant traction in the Ethereum ecosystem due to their ability to reduce gas fees and increase transaction throughput while maintaining security. For example, Arbitrum[arbtirumArbitrumScalable] and Optimism[optimismHome] are leading platforms that use Optimistic Rollups, while zkSync[zksyncZKsync] and StarkWare[starkwareHomepage] leverage ZK-Rollups.
@@
State channels allow participants to conduct multiple off-chain transactions, with only the final state being committed to the blockchain[McCorry2019YouSM]. This method is highly efficient for repeated interactions, such as payments or gaming applications, where the parties involved can settle their accounts off-chain and only submit the final balance to Ethereum.
@@
Plasma is a framework that allows for scalable and secure off-chain transactions[Poon2017PlasmaS]. Plasma chains are separate blockchains that run alongside the Ethereum main chain. They handle large volumes of transactions off-chain and interact with the main chain for security and settlement. Plasma chains can process thousands of transactions per second while ensuring that users can always retrieve their funds from the main chain if something goes wrong with the Plasma chain.
Sidechains[Singh2020SidechainTI], like Plasma, operate independently of the Ethereum main chain but still interact with it for finality and security. Sidechains have their own consensus mechanisms and can offer faster transaction times and lower fees. A popular example of a sidechain is Polygon (formerly Matic), which has become one of the leading Layer 2 solutions by providing fast and low-cost transactions.
@@ 12.3.3 Network Load from Decentralized Applications (dApps)
dApps can suffer from the same scalability issues as the Ethereum network itself, including high gas fees and slow transaction times. Popular dApps, particularly those in the DeFi space, often experience congestion during peak periods, making it difficult for users to interact with them affordably. For instance, during the DeFi summer of 2020, gas fees skyrocketed as users rushed to participate in yield farming[Cousaert2021SoKYA], making smaller transactions prohibitively expensive.
@@
High-frequency activities, such as algorithmic trading on decentralized exchanges (DEXs) or frequent NFT minting, can place additional strain on the network[Brolley2022OnDemandFT]. These types of activities require a high number of transactions in a short period, leading to network congestion and increased gas prices.
@@
To overcome scalability challenges, many dApps are beginning to integrate Layer 2 solutions such as Rollups, state channels, and Plasma. For example, Uniswap[uniswapUniswapInterface], a leading decentralized exchange, has integrated with Optimistic Rollups to offer faster and cheaper trades. Similarly, gaming dApps are using state channels to allow users to interact with the platform off-chain, reducing the cost of frequent in-game transactions.
@@ 13.1.1 Definition of Proof of Stake
Proof of Stake (PoS) is a consensus mechanism that is utilized in blockchain networks to determine how the next block in a blockchain is chosen and validated[Wendl2022TheEI]. Unlike Proof of Work (PoW), where participants (miners) solve complex cryptographic puzzles to create new blocks, PoS selects validators based on the amount of cryptocurrency they have ”staked” or locked up in the network[Trivedi2023EnhancedSF].
In PoS, the idea is that those who have a larger stake in the network have more to lose if the system is compromised[Trivedi2023EnhancedSF]. Hence, validators are selected to propose and validate new blocks in a way proportional to the amount of cryptocurrency they hold and are willing to stake. Validators are chosen through a pseudo-random selection process, with factors such as the size of the stake and the validator’s age (the time they have been staking) playing a role.
This method eliminates the need for intensive computational work, which is one of the main criticisms of PoW[Miraz2021EvaluationOG]. Instead, PoS is considered energy-efficient, as it removes the requirement for high energy consumption seen in mining operations.